what is.
Dave is 4 years older than Colin. Ruth is three times as old as Dave. The total of their three ages is 56 Work out Ruth's age

Answer by Alan3354(69443)

Dave is 4 years older than Colin. Ruth is three times as old as Dave. The total of their three ages is 56 Work out Ruth's age
-----------
D = C + 4
R = 3D = 3(C+4) = 3C+12
----
C + (C+4) + (3C+12) = 56
5C = 40
C = 8
---
R = 3C+12 = 36